# LightMem: Lightweight and Efficient Memory-Augmented Generation

### v1

@@ Abstract
Despite their remarkable capabilities, Large Language Models (LLMs) struggle to effectively leverage historical interaction information in dynamic and complex environments. Memory systems enable LLMs to move beyond stateless interactions by introducing persistent information storage, retrieval, and utilization mechanisms. However, existing memory systems often introduce substantial time and computational overhead. To this end, we introduce a new memory system called LightMem, which strikes a balance between the performance and efficiency of memory systems. Inspired by the Atkinson–Shiffrin model of human memory, LightMem organizes memory into three complementary stages. First, cognition-inspired sensory memory rapidly filters irrelevant information through lightweight compression and groups information according to their topics. Next, topic-aware short-term memory consolidates these topic-based groups, organizing and summarizing content for more structured access. Finally, long-term memory with sleep-time update employs an offline procedure that decouples consolidation from online inference. Experiments on LongMemEval with GPT and Qwen backbones show that LightMem outperforms strong baselines in accuracy (up to 10.9% gains) while reducing token usage by up to 117×, API calls by up to 159×, and runtime by over 12×. The code is available at https://github.com/zjunlp/LightMem.
@@ 1 Introduction
In contrast, human memory offers a compelling model of efficiency and adaptability (Liu et al., 2025a). It operates through a hierarchical architecture: sensory memory pre-filters incoming stimuli before they reach short-term memory, which actively integrates, manipulates, and reasons over task-relevant content. Over time, salient information is selectively consolidated into long-term memory processing in sleep time. This dynamic, multi-stage system balances retention, compression, and retrieval, striking an optimal trade-off between performance and resource usage.
@@ 2.1 Conventional Memory Systems for LLMs
Although implementation details and order vary across methods, LLM memory systems can often be formalized as a three-stage model for LLMs: (I) Raw data $D$ are first processed at a chosen level of granularity, $D^{(g)}=f_{\text{gran}}(D;g),g\in\{\text{turn},\text{session},\text{topic}\}$, so that information can be structured to reflect different contextual scopes. (II) The segmented data $D^{(g)}$ are subsequently summarized or extracted to generate memory units, $U=f_{\text{sum}}(D^{(g)})$, which are then stored and organized within structural backends such as vector databases or knowledge graphs, enabling long-term retention. (III) Many systems incorporate an updating mechanism to mitigate issues such as context conflicts or outdated information, $M^{\prime}=f_{\text{update}}(M,R;U)$, where $M$ denotes the existing memory bank, $R$ represents the newly generated memory units, and $U$ specifies the update or forgetting policy.
@@ 2.2 Atkinson–Shiffrin Human Memory Model
Figure 2: Human memory system.
Following the Atkinson–Shiffrin human memory model (Atkinson & Shiffrin, 1968), raw environmental information in human brain is first briefly retained in sensory memory as shown in Figure 2, which enables rapid pre-attentive feature extraction and filtering, effectively serving as a form of pre-compression. The processed output can then enter short-term memory (STM), where information and interaction sequences are preserved for tens of seconds to minutes, supporting secondary filtering and more deliberate processing. In contrast, long-term memory (LTM) provides durable storage and undergoes continuous reorganization through updating, abstraction, and forgetting. Importantly, Rasch & Born (2013) highlight that sleep plays a critical role in this reorganization, as oscillatory activity during sleep facilitates the integration and consolidation of memory systems.
@@ 2.3 Limitations of Existing LLM Memory Systems
Figure 3: The LightMem architecture. Our LightMem consists of three modules: a) implements an efficient Sensory Memory Module that selectively preserves salient information from raw input, b) realizes a concise STM Module for transient information processing, and c) provides an LTM module designed to minimize retrieval latency.
@@
Analogous to the human memory, we design LightMem as shown in Figure 3, which consists of three light modules: Light1 implements an efficient Sensory Memory Module that selectively preserves salient information from raw input, Light2 realizes a concise STM Module for transient information processing, and Light3 provides an LTM module designed to minimize retrieval latency.
@@ 3.1 Light1: Cognitive-inspired sensory memory
Specifically, dialogue scenarios possess natural semantic units, namely the conversational turn. We construct a turn-level attention matrix $M\in\mathbb{R}^{n\times n}$. $\mathcal{B}_{1}$ are identified as local maxima in the sequence $\{M_{k,k-1}\}$, i.e., the sub-diagonal elements of $M$ corresponding to attention between consecutive sentences. The detailed process of $\mathcal{B}_{1}$ and illustrative cases are provided in Appendix B.1. To mitigate attention sinks and dilution in attention-based methods, we compute semantic similarity between adjacent turns near each candidate boundary in $\mathcal{B}_{1}$. Boundaries with similarity below threshold $\tau$ form set $\mathcal{B}_{2}$, which helps determine the final topic boundaries $\mathcal{B}$.
@@ 3.3 Light3: Long-term Memory with Sleep-time Update
Table 1: Effectiveness and efficiency comparison. The token usage is in thousands. – indicates no value for the metric. Bold denotes the best result, underline the second-best. $r$ denotes the compression rate. thth denotes the capacity threshold of the STM buffer, measured in tokens. Each pair of $r$ and thth corresponds to two rows: one for online soft update and one for offline update. OP-update denotes the offline parallel update process of LightMem.
@@
Experimental Details. (1) Our experiments adopt a realistic Incremental Dialogue Turn Feeding setting, where dialogue turns are processed sequentially as they arrive. This mirrors practical applications where the full dialogue history is unavailable until the conversation concludes (Hu et al., 2025). (2) For considerations of both efficiency and effectiveness, we employ LLMLingua-2 as our pre-compressor throughout all subsequent experiments. (3) The attention scores for topic segmentation are also obtained using LLMLingua-2; therefore, the size of the sensory memory buffer matches the model’s context window length, which is 512 tokens.
@@
Dataset & Baseline Methods. We use the well-known dataset LongMemEval (Wu et al., 2025) to evaluate memory ability. Specifically, we adopt LongMemEval-S, which consists of 500 dialogue histories, each containing an average of 50 sessions and 110k tokens. Under the evaluation protocol, we choose it because of its balance between dialogue length and computational cost feasibility. We compare LightMem against several representative baselines of conversational memory modeling. ① Full Text, ② Naive RAG, ③ LangMem (LangChain, 2025), ④ A-MEM (Xu et al., 2025), ⑤ MemoryOS (Kang et al., 2025), ⑥ Mem0 (Chhikara et al., 2025). In addition, all methods use GPT-4o-mini and Qwen3-30B-A3B-Instruct-2507 as the LLM backbones. Details on dataset, baselines, and experimental settings are provided in the Appendix C.
@@
Metrics. We evaluate these methods using both effectiveness and efficiency metrics. For effectiveness, we report Accuracy (ACC), defined as the proportion of correctly answered questions. Following prior work, evaluation is conducted with GPT-4o-mini as an LLM judge, guided by a detailed evaluation prompt (see Appendix D.1). For efficiency, we focus on tracking the computational costs of LLM invocations, all averaged across the entire dataset. We define two key operations of memory management: the Summary process as memory distillation and summarization operations, and the Update process as memory consolidation or deletion operations. For both processes, we report the token consumption from LLM calls, including input tokens, output tokens, and total token usage (in thousands). Additionally, we track API Calls counting the total number of LLM invocations, and Runtime recording the execution time in seconds. Our focus is on the overhead of the aforementioned processes. The costs associated with subsequent retrieval and QA are not specifically analyzed, due to their minimal and comparable token/time overhead.
Table 2: The impact of LightMem compression ratio $r$ and STM buffer threshold $th$ is reported here. Due to space limitations, we only present a subset of representative results of the online soft update results, with more results provided in the Appendix 4.
Model thth $r$ ACC Input (k) Output (k) Total (k) Calls Time
GPT
256 0.5 64.29 20.80 10.01 30.81 25.67 302.69
256 0.6 67.68 24.58 10.53 35.11 30.47 329.61
256 0.7 65.68 27.66 9.97 37.63 34.26 403.59
512 0.6 63.74 16.23 9.45 25.68 15.63 266.98
512 0.7 68.64 18.88 9.37 28.25 18.43 283.76
512 0.8 66.67 21.55 8.59 30.14 21.11 268.97
1024 0.6 59.68 10.34 7.68 18.20 7.69 177.45
1024 0.7 64.68 12.93 6.90 19.83 8.25 209.12
1024 0.8 64.35 14.86 6.28 21.14 9.43 216.08
Qwen
512 0.4 58.57 11.03 17.00 28.03 10.11 421.74
512 0.6 66.57 16.22 19.50 35.72 15.40 471.09
512 0.8 67.37 21.35 19.36 40.71 20.98 461.02
768 0.4 61.95 9.01 16.14 25.15 6.54 357.13
768 0.6 73.20 13.19 19.21 32.40 9.97 417.13
768 0.8 64.95 16.94 19.06 36.00 13.09 420.14
1024 0.4 53.91 8.02 15.44 23.46 4.83 300.56
1024 0.6 65.67 11.50 18.21 29.71 7.18 396.35
1024 0.8 68.69 14.82 18.49 33.31 9.43 355.71
@@
As shown in Table 1, LightMem demonstrates superior effectiveness and efficiency across three parameter settings on both GPT and Qwen backbones. We evaluated LightMem in two stages: one before the sleep time update, i.e., at the end of the online soft update phase when the memory construction is completed; and the other after the offline update.
@@
Across all three parameter settings, LightMem achieves superior performance on nearly all metrics and both LLM backbones, demonstrating its robustness and high flexibility.
@@ 4.3 Analysis of Pre-Compressing Submodule
Impact of $r$ on Performance. As shown in Tables 2 and  4, The optimal $r$ for ACC is dependent on the STM buffer threshold $th$. For smaller thresholds ($th\in\{0,256\}$), an $r$ of 0.6 achieves the highest ACC. In contrast, for larger thresholds ($th\in\{512,1024\}$), a higher retention rate of $r=0.7$ performs best. This suggests greater buffer capacity enables effective use of richer, less-compressed information, leveraging LLMs’ advanced long-context processing to mitigate the “lost in the middle” phenomenon. On average, the optimal $r$ for ACC is 0.6, reflecting a trade-off between information compression rate and the quantity of information in the STM buffer. In terms of efficiency, a lower $r$ generally leads to higher efficiency, as it triggers the buffer threshold less frequently under the same $th$, resulting in fewer API calls and lower token consumption.
@@ 4.6 Analysis of Sleep-Time Update
6 conclusion and Future Work
In this work, we introduced LightMem, a lightweight and efficient memory framework designed to address the significant overhead of memory systems for LLM agents. Inspired by the multi-stage Atkinson-Shiffrin human memory model, LightMem’s architecture effectively filters, organizes, and consolidates information. Our empirical evaluation demonstrates that this approach maintains strong task performance while sharply reducing computational costs.
Offline Update Acceleration. We plan to enhance the efficiency of LightMem’s update phase by incorporating pre-computed key–value (KV) caches. This optimization allows KV cache pre-computation to be performed offline, significantly accelerating memory consolidation and reducing runtime overhead during interactive sessions.
Knowledge Graph-based Memory. To better address complex multi-hop reasoning challenges, we aim to integrate a lightweight knowledge graph-based memory module. This module will support explicit relational reasoning and structured information retrieval, enabling the model to perform more interpretable and compositional reasoning across interconnected knowledge entities.
Multimodal Memory Extension. We further plan to develop a multimodal memory mechanism that allows LightMem to adapt to multimodal models and scenarios. Such an extension is crucial for embodied agents and real-world applications where visual, auditory, and textual information jointly contribute to memory formation and retrieval.
Parametric–Nonparametric Synergy. We will explore the collaborative mechanisms between parametric and non-parametric memory components. To bridge the gap between the two paradigms, we aim to enable more flexible, synergistic knowledge utilization, combining the efficiency of parametric representations with the interpretability and adaptability of non-parametric storage.
@@
To ensure the reproducibility of this work, we introduce the detailed implementations for LightMem are provided in in Section 3, Appendix B. Additionally, we plan to release our source code in the future to further support reproducibility. These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
Appendix A Usage of LLMs

### v3

@@ Abstract
[…]
@@ 1 Introduction
In contrast, human memory efficiently processes information through a hierarchical system: sensory memory pre-filters stimuli, short-term memory actively integrates and reasons over relevant content, and long-term memory selectively consolidates salient information in sleep time.
@@ 2.1 Conventional Memory Systems for LLMs
We describe mainstream memory architectures pipeline in terms of two major stages. (I) Memory Bank Construction. This stage can be further decomposed into three sub-stages: (a) Raw data $D$ are first processed at a chosen level of granularity, $D^{(g)}=f_{\text{seg}}(D;g),g\in\{\text{turn},\text{session},\text{topic}\}$ in dialog scenario; (b) The segmented data $D^{(g)}$ are then summarized or extracted to generate memory entries, $E=f_{\text{sum}}(D^{(g)})$, which are stored and organized within structural backends such as vector databases or knowledge graphs to enable long-term retention; (c) Many systems incorporate an updating mechanism to mitigate issues such as context conflicts or outdated information, $M^{\prime}=f_{\text{update}}(M,R;U)$, where $M$ denotes the existing memory bank, $R$ represents newly generated memory entries, and $U$ specifies the update or forgetting policy. (II) Retrieval and Usage. When a new user query arrives, the system retrieves relevant entries from the memory bank, integrates them with the query to construct the final prompt, and then invokes the model to produce a response.
@@ 2.2 Atkinson–Shiffrin Human Memory Model
Following the Atkinson–Shiffrin human memory model (Atkinson & Shiffrin, 1968), raw environmental information in human brain is first briefly retained in sensory memory, which enables rapid pre-attentive feature extraction and filtering, effectively serving as a form of pre-compression. The processed input can then enter short-term memory (STM), where information and interaction sequences are preserved for tens of seconds to minutes, supporting secondary filtering and more deliberate processing. In contrast, long-term memory (LTM) provides durable storage and undergoes continuous reorganization through updating, abstraction, and forgetting. Importantly, Rasch & Born (2013) highlight that sleep plays a critical role in this reorganization, as oscillatory activity during sleep facilitates the integration and consolidation of memory systems.
@@ 2.3 Limitations of Existing LLM Memory Systems
Figure 2: The LightMem architecture. LightMem consists of three modules: a) An efficient Sensory Memory Module, b) a topic aware STM Module, and c) an LTM module updated in sleep time.
@@
Analogous to the human memory, we design LightMem as shown in Figure 2, which consists of three light modules: Light1 implements an efficient Sensory Memory Module that selectively preserves salient information from raw input (§3.1), Light2 realizes a topic-aware STM Module for transient information processing (§3.2), and Light3 provides an LTM module designed to minimize test time update latency (§3.3) with a sleep time update mechanism. The overall pipeline framework of LightMem, its specific models, and comparisons with other memory frameworks are presented in Appendix A.1. The complexity analysis for LightMem’s efficiency gains is in Section 4.
@@ 3.1 Light1: Cognitive-inspired sensory memory
Specifically, dialogue scenarios possess natural semantic units, namely the conversational turn. We construct a turn-level attention matrix $M\in\mathbb{R}^{n\times n}$. $\mathcal{B}_{1}$ are identified as local maxima in the sequence $\{M_{k,k-1}\}$, i.e., the sub-diagonal elements of $M$ corresponding to attention between consecutive sentences. The detailed process of $\mathcal{B}_{1}$ and illustrative cases are provided in Appendix C.1. To mitigate attention sinks and dilution in attention-based methods, we compute semantic similarity between adjacent turns near each candidate boundary in $\mathcal{B}_{1}$. Boundaries with similarity below threshold $\tau$ form set $\mathcal{B}_{2}$, which helps determine the final topic boundaries $\mathcal{B}$.
@@ 3.3 Light3: Long-term Memory with Sleep-time Update
4 Complexity analysis about LightMem
Method Summary Tokens Update Tokens API Calls Runtime
Baselines
$N(L_{\text{sum-in}}+T+L_{\text{sum-out}})$ $M_{1}R_{1}(L_{\text{up-in}}+L_{\text{up-out}}) NN$ $O(N)$
LightMem $\frac{Nr^{x}T}{th}(L_{\text{sum-in}}+th+L_{\text{sum-out}})$ $\frac{Nr^{x}T}{th}M_{2}R_{2}(L_{\text{up-in}}+L_{\text{up-out}})$ $\frac{Nr^{x}T}{th}$ $O\Big(\frac{Nr^{x}T}{th}\Big)$
Table 1: Complexity comparison between LightMem and other memory systems. The specific definitions of each symbol are provided in the Appendix A.2.
As shown in Table 1, we consider a dialogue with $N$ turns, each containing on average $T$ tokens. In conventional memory systems, each turn triggers a summarization call, consuming $L_{\text{sum-in}}+T+L_{\text{sum-out}}$ tokens and totaling $N(L_{\text{sum-in}}+T+L_{\text{sum-out}})$ tokens with $N$ API calls. Each summarization produces $M_{1}$ memory entries, a fraction $R_{1}$ of which retrieve at least one relevant neighbor and trigger an update, resulting in an update-token cost of $NM_{1}R_{1}(L_{\text{up-in}}+L_{\text{up-out}})$.
In LightMem, each turn is first passed through iterative pre-compression submodule, retaining only $r^{x}T$ tokens after $x$ iterations, and appended to a short-term memory (STM) buffer of capacity $th$. Summarization is triggered only when the buffer reaches capacity, yielding $\frac{Nr^{x}T}{th}$ summarization calls, each consuming $L_{\text{sum-in}}+th+L_{\text{sum-out}}$ tokens. Each summarization produces $M_{2}$ memory entries, but stricter retrieval constraints, including semantic similarity and timestamp filtering, reduce the fraction $R_{2}$ that trigger updates. Hence, the update phase involves $\frac{Nr^{x}T}{th}M_{2}R_{2}$ calls, with a total token cost of $\frac{Nr^{x}T}{th}M_{2}R_{2}(L_{\text{up-in}}+L_{\text{up-out}})$.
Overall, LightMem requires only $\frac{Nr^{x}T}{th}$ API calls for both summarization operations, substantially reducing token usage and call frequency compared to other systems. Correspondingly, the runtime complexity of other memory systems is $O(N)$, while LightMem achieves a reduced runtime of $O\Big(\frac{Nr^{x}T}{th}\Big)$, reflecting the efficiency gain from compressed summarization and selective updates.
Table 2: Effectiveness and efficiency comparison on LongMemEval-S. The token usage is in thousands. – indicates no value for the metric. Bold denotes the best result, underline the second-best. $r$ denotes the compression rate. thth denotes the capacity threshold of the STM buffer, measured in tokens. Each pair of $r$ and thth corresponds to two rows: one for online soft update and one for offline update. OP-update denotes the offline parallel update process of LightMem.
@@ 4 Complexity analysis about LightMem
Experimental Details. (1) Our experiments adopt a realistic Incremental Dialogue Turn Feeding setting, where the entire dialogue history is fed and processed at the turn level, one turn at a time. This reflects practical scenarios where interactions between user and model is incrementally formed turn by turn. (Hu et al., 2025). (2) For considerations of both efficiency and effectiveness, we employ LLMLingua-2 as our pre-compressor throughout all subsequent experiments. (3) The attention scores for topic segmentation are also obtained using LLMLingua-2, the size of the sensory memory buffer is 512 tokens. All specific models used in this paper, can be found in Table 5.
@@
Datasets & Baseline Methods. We use two well-known datasets, LongMemEval (Wu et al., 2025) (specifically the LongMemEval-S split) and LoCoMo (Maharana et al., 2024) to evaluate memory ability. We compare LightMem against several representative baselines of conversational memory modeling. ① Full Text, ② Naive RAG, ③ LangMem (LangChain, 2025), ④ A-MEM (Xu et al., 2025), ⑤ MemoryOS (Kang et al., 2025), ⑥ Mem0 (Chhikara et al., 2025). In addition, all methods use GPT-4o-mini and Qwen3-30B-A3B-Instruct-2507 as the LLM backbones. Details on dataset, baselines, and experimental settings are provided in the Appendix D.
@@
Metrics. We evaluate these methods using both effectiveness and efficiency metrics. For effectiveness, we report Accuracy (ACC), defined as the proportion of correctly answered questions. The evaluation is conducted with GPT-4o-mini as an LLM judge, guided by a detailed evaluation prompt (see Appendix E.1). For efficiency, we focus on tracking the computational costs of the LLM invocations in memory bank construction stage (see Section 2.1), all averaged across the entire dataset, as it is the one tied to the design and implementation differences of memory systems. The retrieval and usage stage is not our focus, because for fair comparison, The $f_{\text{retrieve}}()$, $f_{\text{chat}}()$ and number of retrieved entries are same among all methods. As a result, their costs exhibit only minor differences, and this stage is largely orthogonal to the design of memory systems, as shown in the table. Within the memory bank construction stage, only the two sub-processes Summary and Update involve the use of LLMs, $f_{\text{sum/extract}}()$ and $f_{\text{update}}()$. So for both processes, we report the token consumption from LLM calls, including input tokens, output tokens, and total token usage (in thousands). Additionally, we track API Calls counting the total number of LLM invocations, and Runtime recording the overall execution time for memory bank construction stage.
Table 3: Effectiveness and efficiency comparison on LoCoMo. Due to space limitations and for ease of comparison, we merge the results before and after LightMem’s offline update into a single row. The ACC reported corresponds to the performance after the offline update.
In Out In Out
[Uncaptioned image] GPT-4o-mini
FullText 71.83
NaiveRAG 63.64
LangMem 57.20 898.27 111.95 1010.22 920.62 2229.37
A-MEM 64.16 182.74 49.29 729.89 187.52 1149.43 1175.47 6060.73
MemoryOS(locomo) [^1] MemoryOS(locomo) is the LoCoMo reproduction script in the MemoryOS library, simplifying the standard version, shown as MemoryOS(regular). 58.25 110.98 33.40 78.08 64.54 287.00 553.45 2422.05
MemoryOS(regular) 54.87 226.86 46.61 177.66 75.34 526.48 1016.06 3332.59
Mem0 61.69 851.32 20.53 632.12 189.42 1693.39 1602.20 4432.87
LightMem(0.7,512) 71.95 73.19 20.13 6.05 0.40 99.76 41.65 848.49
LightMem(0.7,768) 70.26 57.54 18.92 3.79 0.23 80.48 29.55 737.80
LightMem(0.8,768) 72.99 62.82 17.95 4.14 0.28 85.19 29.83 815.32
[Uncaptioned image] Qwen3-30B-A3B-Instruct-2507
FullText 74.87
NaiveRAG 66.95
LangMem 60.53 1004.35 138.02 1142.37 1005.37 2268.57
A-MEM 56.10 158.29 60.85 924.19 483.51 1626.80 1175.40 5543.90
MemoryOS(locomo) 61.04 122.21 53.12 104.43 81.75 361.51 414.70 1269.70
MemoryOS(regular) 51.30 228.85 51.60 242.27 143.63 666.35 1004.60 1982.20
Mem0 43.31 827.09 18.64 763.88 189.80 1799.40 1614.50 4540.70
LightMem(0.6,768) 71.36 56.68 34.14 8.31 0.74 99.87 29.10 815.70
LightMem(0.8,1024) 72.60 61.38 36.33 9.86 0.88 108.45 32.00 1079.40
@@
As shown in Table 2 and Table 3, LightMem demonstrates superior effectiveness and efficiency on both datasets across both GPT and Qwen backbones. For a fair comparison, all efficiency metrics for LightMem in the following analysis refer to the combined online and offline costs.
@@
[…]
LightMem achieves superior performance on nearly all metrics and both LLM backbones, while demonstrating robust performance and efficiency on both LongMemEval and LoCoMo, highlighting its generalizability across different models and scenarios.
@@ 5.3 Analysis of Pre-Compressing Submodule
Impact of $r$ on Performance. As shown in Tables 8 and  9, The optimal $r$ for ACC is dependent on the STM buffer threshold $th$. For smaller thresholds ($th\in\{0,256\}$), an $r$ of 0.6 achieves the highest ACC. In contrast, for larger thresholds ($th\in\{512,1024\}$), a higher retention rate of $r=0.7$ performs best. This suggests greater buffer capacity enables effective use of richer, less-compressed information, leveraging LLMs’ advanced long-context processing to mitigate the “lost in the middle” phenomenon. On average, the optimal $r$ for ACC is 0.6, reflecting a trade-off between information compression rate and the quantity of information in the STM buffer. In terms of efficiency, a lower $r$ generally leads to higher efficiency, as it triggers the buffer threshold less frequently under the same $th$, resulting in fewer API calls and lower token consumption.
@@ 5.6 Analysis of Sleep-Time Update
7 conclusion
In this work, we introduced LightMem, a lightweight and efficient memory framework designed to address the significant overhead of memory systems for LLM agents. Inspired by the multi-stage Atkinson-Shiffrin human memory model, LightMem’s architecture effectively filters, organizes, and consolidates information. Our empirical evaluation demonstrates that this approach maintains strong task performance while sharply reducing computational costs. In the near future, we plan to accelerate LightMem’s update phase via offline pre-computed KV caches, reducing runtime overhead. We aim to integrate a lightweight knowledge graph memory for explicit multi-hop reasoning and structured retrieval. A multimodal memory extension will enable adaptation to visual, auditory, and textual inputs in embodied and real-world scenarios.
@@
To ensure the reproducibility of this work, we introduce the detailed implementations for LightMem are provided in in Section 3, Appendix C. Additionally, we plan to release our source code in the future to further support reproducibility. These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
Appendix A Background Details
A.1 background about current memory systems
We describe both the mainstream memory architectures and the LightMem pipeline in terms of two major stages. The first is the memory bank construction stage, which can be further decomposed into the three sub-stages (I), (II), and (III) described in the Section 2.1. The second major stage concerns the usage of the memory system, which consists of retrieval and question answering (QA).
Memory Bank Construction
As shown in Table 4, we detail the workflows of the three sub-stages (I), (II), and (III) for naive RAG, prevailing memory systems, and our LightMem. It can be observed that baseline memory systems typically perform their update stage during user–model interaction, which introduces substantial test-time latency. In contrast, LightMem decouples this update process from online interaction, thereby significantly reducing test-time latency. All models involved in these processes are listed in Table 5. As shown, LightMem introduces only one additional model, LLMlingua-2,beyond those used by baseline methods. This model follows a lightweight BERT architecture and requires less than 2GB of GPU memory during inference, rendering its overhead negligible. Moreover, for fairness, the latency introduced by this component is fully accounted for in our reported Runtime metric.
Table 4: The mainstream memory architectures and the LightMem pipeline of memory bank construction stage. Black-font processes denote those executed during online test-time interactions, whereas red-font processes denote those executed offline.
Method (I) Segment (II) Summary/Extrct (III) Update
NaiveRAG
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{index}}()\to\{\text{emb}_{i}\}$ \
Other
Memory
Systems
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{sum/extract}}()\to\{\text{memory entry}_{i}\}$
$\to f_{\text{index}}()\to\{\text{emb}_{i}\}$
$\to f_{\text{retrieve}}()\to\{\text{related entry}_{i}\}$
$\to f_{\text{update}}()$
$\to\{\text{add, delete, update, merge...}\}$
LightMem
$\text{Raw dialog}\to f_{\text{seg}}()$
$\to\{\text{seg}_{i}\}$
$\to f_{\text{pre\_compress}}()$
$\to\{\text{comp\_seg}_{i}\}$
$\to\text{sensory buffer full}\to$
$f_{\text{topic}}()\to$
$\{\text{topic-wise comp\_seg}_{i}\}$
$\to\text{STM buffer full}$ $\to f_{\text{sum/extract}}()$
$\to\{\text{topic}_{i},\{\text{memory entry}_{j}\}\}$
$\to f_{\text{index}}()\to\{\text{topic}_{i},\{\text{emb}_{j}\}\}$
Offline update trigger
$\{\text{every entry}_{i}\}\to f_{\text{retrieve}}()$
$\to\{\text{related entry}_{j}\}\to\{\text{update queue}\}$
All update queues established
$\to\text{parallel }f_{\text{update}}()$
$\to\{\text{add, delete, update, merge...}\}$
Function Model / Strategy Implementation in This Paper
$f_{\text{seg}}()$ Segmentation strategy Turn-level granularity input
$f_{\text{index}}()$ Embedding model all-MiniLM-L6-v2
$f_{\text{sum/extract}}()$ System backbone model GPT-4o-mini; Qwen3-30B-A3B-Instruct-2507
$f_{\text{retrieve}}()$ Retrieval strategy Cosine similarity vector retrieval
$f_{\text{update}}()$ System backbone model GPT-4o-mini; Qwen3-30B-A3B-Instruct-2507
$f_{\text{pre\_compress}}()$ Token compression model LLMlingua-2
$f_{\text{topic}}()$ Topic segmentation model LLMlingua-2
$f_{\text{chat}}()$ Chat model GPT-4o-mini; Qwen3-30B-A3B-Instruct-2507
Table 5: Mapping between functions, their roles, and the concrete models used in this paper. Black-font entries denote models shared by both LightMem and baseline methods, whereas red-font entries denote models unique to LightMem.
Retrieval and Usage
After the memory bank construction stage, we obtain an up-to-date memory bank. When a new user query arrives, the memory system use $f_{\text{retrieve}}()$ to retrieve relevant entries from this repository, appends them to the query, and then prompts the chat model $f_{\text{chat}}()$ to produce a response.
A.2 Notation and Complexity Details
Table 6: Notation used in complexity analysis (§Section 4).
Symbol Definition
NN Total number of turns in a dialogue history.
$T$ Average number of tokens per turn.
$r$ Token compression rate (as defined in the main paper). After one compression step, only a fraction $r$ of tokens is retained.
$x$ Number of compression iterations. In LightMem, the pre-compress module may be invoked multiple times for the same message to remove redundancy until the message is sufficiently compact. This occurs frequently in datasets such as LongMemEval. All time costs are included in runtime metrics.
thth Capacity of the Short-Term Memory (STM) buffer, as defined in the paper.
$L_{\text{sum-in}}$ / $L_{\text{sum-out}}$ Number of tokens in the input prompt template and output of a single backbone LLM call for summarization. These are similar across memory frameworks.
$M_{1}$ / $M_{2}$ Number of memory entries produced from a single summarization operation under Other Memory Systems ($M_{1}$) and LightMem ($M_{2}$).
$L_{\text{up-in}}$ / $L_{\text{up-out}}$ Number of tokens in the input prompt template and output of a single backbone LLM call for memory update. Similar across frameworks.
$R_{1}$ / $R_{2}$ Proportion of summary entries that successfully retrieve at least one relevant memory entry (triggering an update) for Other Memory Systems ($R_{1}$) and LightMem ($R_{2}$). Some entries do not retrieve any relevant counterparts and thus do not trigger updates.
Appendix B Usage of LLMs